When the sum of two consecutive odd numbers in increased by 15, the answer is 59. What are the two odd consecutive numbers?
Answer by jcandy2011(2)

The two odd consecutive numbers are 21 & 23
(X+Y) + 15 = 59
(X+Y) = 59-15
(X+Y) = 44
If the numbers have to be two consecutive odd numbers, they have to 21 & 23 only.